SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the March 30 Background Remediation Steering Committee meeting

Presents : P. Burchat, T Geld, G. Dubois-Felsmann, B. Jacobsen, T. Mattison, W. Kozanecki, G. Wormser

  1. Background group organization
  2. The committee endorsed the proposed structure. The need for a dedicated liaison group hypernews group and web page was expressed. (action item : create such things! GW, TG). The steering committee will stay in place

    in the new unified background group.

  3. Start up strategies
  4. The committee expressed the view that discussions including BABAR , PEP-II and the BR representatives should start rapidly to converge on the guidelines for the may/august run starting from the physics goals (Stew et al.)

  5. Liaison group
  6. A message has to be sent to BABAR management to remind them that it is now pressing to announce to the PIs that liaison shifts are integral part of BABAR shift organisation , that they will be counted therefore as others shifts, that they are very important, etc..(GW, D. Hitlin)

    To increase the BABAR knowledge on the liasion shifts, they will be asked to follow the regular BABAR shift training

  7. Epics and sensors readiness
  8. T. Geld and A. McKemey preparing the BABAR detailed sensors table, with status information. The complete list with full epics name will be sent to Jim Olsen who has kindly volunteered to prepare the global BABAR-PEP-II summary page.

    Two main issues were discussed:

    -Availability in EPICS of OEP data. There is code available from the EPICS collaboration to generate EPICS Channel Access (CA) data from a UNIX process, and that some BaBar people have played with that code and shown that it works. It has not been seriously imported into the BaBar SRT (Software Release Tools) environment. More importantly, though, there is no code to connect DHP histograms and scalers to this Unix CA server. That's a substantial project, and its priority has to be evaluated against others.

    Action item : discuss with Online management this issue (GW, A. Lankford)

    -How to take data just after injection when babar is not in a physics data taking configuration. Gregory mentions that the simplest way is to define another configuration (e.g. ,BABAR_Tuning) to get the runnable flag on even with voltages etc different from nominal settings. One key issue is to decide how to revise the software interlocks that are supposed to prevent any data-taking when PEP-II doesn't have stable beams. An additional "state" of PEP-II (beam tuning) may have to be invented.

    The precise mechanism to generate and use these configurations has to be discussed in a subcomittee including Gerry Abrams. (GW, G. Abrams)(Note that these configurations will also include the trigger configuration that could be very different from the normal BABAR trigger, for the purpose of background assessment (for instance 2 Hz random, 5 Hz tight trigger, 5 Hz prescaled normal trigger)

  9. Detailed start up check list
  10. This is not felt as being urgent. Scheduling negociations with babar and pep-ii for getting the dedicated time to set up and test the protection system, the sensors, the background status, etc will be made during the coming months. Tom emphasizes that much of this commissioning can be done early even with no beam.

  11. PRV0 and background simulation
  12. Despite that the fact that a big flaw has been found in the simulation (the Q2 magnet was filled with air), the committee recommends to pursue vigorously the study of PRv0 background frames since they represent a background level apparently close to the tolerable level. To that effect, the subdetectors that did not present their initial results on PRV0, and especially the EMC because of potentially severe background effects, will be requested to do so in an upcoming BR video meeting,(GW, H. Marsiske,..) well before the software week.

    Meanwhile, work is of course going on to fix the simulation and regenerate some frames.

  13. Software week preparation:

The work described above and discussions on strategy issues. They should be sufficiently advanced so that a report is made at the software week.